# R3-RAG: Learning Step-by-Step Reasoning and Retrieval for LLMs via Reinforcement Learning

### v1

@@ 3.3 Reinforcement Learning
Although the cold start can let the model learn the manner of iterative reasoning and retrieval, it is hard to make the model fully explore the external retrieval environment and thus makes the model limited by the cold-start data. To address these challenges, we propose to use RL to further harness the model’s own abilities to better reason and retrieve external information. Specifically, we propose two rewards for R3-RAG: 1) answer correctness for outcom rewards, which judge whether the trajectory leads to correct answers by ground truth; 2) relevance-based document verification for process reward, encouraging the model to retrieve documents that are highly relevant to user questions. In this section, we first introduce the reward design and then the training objective.
@@
The high-quality outcome reward is critical for the model to generate the correct trajectory  (DeepSeek-AI et al., 2025). For question-answering, string matching is usually used to judge whether the prediction is correct according to the gold answer. However, since the same entity can have different variants, the match-based judgment can be inaccurate, which introduces extra noise to the training signal. In this paper, we propose to combine the match-based judgment and model-based judgment as follows:
@@
where $a$ is the predicted answer, and $\mathrm{Acc}_{match}$ and $\mathrm{Acc}_{model}$ are match-based and model-based answer judgement.
@@
Using only outcome reward is challenging because it provides limited fine-grained training signal and results in sparse rewards, which degrades the effect of RL. We further propose relevance-based document verification to give fine-grained guidance to the process of iterative retrieval. In each retrieval step, we calculate the document relevance as:
$Rel(d_{i}^{j})=\mathrm{LLM}(I_{relevance},q_{i},d_{i}^{j}),$ (3)
where $I_{relevance}$ is the instruction for the LLM to judge the relevance between question $q_{i}$ and documents $d_{i}^{j}$, with scores ranging from $0,0.1,0.2,\dots$ to $1$. Since $q_{i}$ and $d_{i}^{j}$ are concatenated and then sent to the LLM, the relevance judgment can take advantage of full token-level interaction and thus guide the model to better retrieve relevant information. Additionally, since retrieving relevant documents in one retrieval step is easier than getting the correct answer in the final step, this relevance-based reward can provide more fine-grained guidance and make it easier for the model to learn how to retrieve comprehensive information.
@@
We also implement a format reward $r_{format}(s_{i}^{j})$ to encourage the model to generate the retrieval invoking in the required format.
@@
The overall reward is the sum of the answer correctness reward, document relevance reward, and format reward. In this way, we can let the model learn how to iteratively invoke retrievers and obtain external information to achieve the final correct answer. To enhance the effect of answer correctness based reward and prevent the training progress from being distracted by other rewards, we use the following factor to adjust the process reward throughout the trajectory:
$r^{*}_{\text{process}}(s_{i}^{j})=r_{\text{process}}(s_{i}^{j})\cdot\lambda_{T}(T_{i}),$ (4)
where $\lambda_{T}(T_{i})$ is determined based on the trajectory outcome. For trajectories with correct answers, $\lambda_{T}(T_{i})$ is positive to reinforce successful reasoning; for incorrect answers or failures to produce answers due to overthinking, it is negative to discourage these behaviors. For trajectories with format errors, $\lambda_{T}(T_{i})$ equals 1 to isolate format penalties to specific steps.
3.3.1 Training Objective
We apply the Proximal Policy Optimization (PPO) on the cold-start model. Guided by rewards from both the reasoning process and the final outcome, we update the model weights according to the following loss function:
$\mathcal{L}_{RL}=$ $\mathbb{E}{(x,y)\sim\mathcal{D}_{RL\_train}}\Big{[}\min\Big{(}\rho(\theta)A(x,y),$
$\text{clip}(\rho(\theta),1-\epsilon,1+\epsilon)A(x,y)\Big{)}\Big{]}$ (5)
$+\beta\mathcal{L}_{KL},$
where $\rho(\theta)=\frac{P(y|x;\theta)}{P(y|x;\theta_{old})}$ represents the probability ratio between the current policy and the old policy, $A(x,y)$ denotes the advantage function estimated using Generalized Advantage Estimation (GAE), $\epsilon$ is the clipping parameter that constrains policy updates, $\mathcal{L}_{KL}$ represents the Kullback-Leibler divergence between the current policy and the reference model to prevent excessive deviation, and $\beta$ controls the strength of the KL penalty (Schulman et al., 2017). Details are provided in Appendix D.
4 Experiments
@@
Table 1: Performance comparison of different models and retrieval methods across multi-hop QA datasets. Our R3-RAG method consistently outperforms all baselines across accuracy (ACC), F1, and exact match (EM) metrics regardless of the base model or retriever used. Bold values indicate best performance in each column and model group. Missing values (-) indicate results not reported in original papers.
@@ 4.1 Datasets and Metrics
We follow Song et al. (2025) to conduct comparison on three multi-hop QA datasets: HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022), using their full dev sets. We evaluate using three complementary metrics: Exact Match (EM) measures strict string match with golden answers; F1 Score quantifies partial matching through token overlap between predicted and golden answers; and Accuracy (Acc) is based on LLM judgment of semantic correctness. For EM and F1, we first extract the short answer from the model’s response and then compare it with the gold answer.
@@ 4.2 Baselines
We compare against several competitive baselines: (1) CoT: the LLMs directly answer the question with Chain-of-Thought; (2) RAG with CoT: the LLMs answer the question based on the documents retrieved by original questions as queries; (3) Iterative retrieval methods: models are prompted to invoke retrieval tools during reasoning, including ReAct Yao et al. (2023), Flare Jiang et al. (2023), Self-ask Press et al. (2023), ITER-RETGEN Shao et al. (2023), and IRCoT Trivedi et al. (2023); and (4) Fine-tuned methods: including Auto-RAG Yu et al. (2024), DeepRAG Guan et al. (2025), and ReSearch Chen et al. (2025).
@@ 4.3 Implementation Details
For the cold-start stage, we create 51,254 solution trajectories from the training sets of HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022). For the reinforcement learning stage, we sample 8,192 instances from the remaining HotpotQA training data. We employ E5-base-v2 (Wang et al., 2022) as our default retriever. During evaluation, the models receive complete retrieval documents, limited to the top-5 per iteration with a maximum of 5 iterations. Unlike prior work  (Zheng et al., 2025), which extracts information from retrieval documents to reduce context length, our approach directly incorporates complete retrieval documents. All methods are evaluated using the same metrics and hyper-parameters. Further implementation details are provided in Appendix E.
@@ 4.4 Main Results
We show the evaluation results of R3-RAG using Llama-3.1-8B and Qwen2.5-7B in Table 1. We see that R3-RAG significantly outperforms baselines on three datasets across different metrics, which shows its best overall reasoning and retrieval performance. Specifically, R3-RAG outperforms the strongest iterative RAG workflow method, IRCOT, by around 15 points on average, which shows that R3-RAG can help the model better learn how to iteratively reason and retrieve step-by-step and thus get more accurate answers. Additionally, R3-RAG also outperforms ReSearch (Chen et al., 2025), which uses RL with outcome-based reward to augment the model’s iterative retrieval ability. This shows the advantages of R3-RAG’s reward design, which not only teaches the model how to get the final answers, but also guides the model to retrieve documents relevant to the question. Meanwhile, R3-RAG achieves significant improvement based on both Qwen2.5-7B and Llama-3.1-8B, and this shows the robustness of R3-RAG. Although R3-RAG’s RL is only trained on the HotpotQA dataset, R3-RAG outperforms the cold-start model on all three datasets, which shows the proposed RL algorithm can generalize well across various datasets and further demonstrates R3-RAG’s transferability for different downstream scenarios.
@@ 5.1 Effect of Different Reward
Methods 2Wiki MuSiQue Avg
@@
Table 2 presents the comparative performance of different reinforcement learning configurations in R3-RAG. Our full model employs both process and outcome rewards. When we remove the process reward component from our complete R3-RAG framework, we observe a performance decline of 1.4% on average (from 46.6% to 45.2%), which shows that fine-grained process rewards effectively guide the model to retrieve more relevant documents at each reasoning step, leading to improved overall performance. Further removing the outcome reward component (thus eliminating all reinforcement learning signals) results in an additional performance drop of 3.6% on average (from 45.2% to 41.6%). These results demonstrate the effectiveness of both reward components. Outcome rewards guide the model to optimize for correct final answers, while process rewards provide granular feedback on the quality of each retrieval operation. By jointly optimizing for both the ultimate goal (correct answers) and the intermediate steps (relevant document retrieval), our approach achieves a more effective reasoning-retrieval strategy.
Figure 3: Impact of the maximum number of iteration steps on HotpotQA and 2WikiMultiHopQA. Results are shown for both Qwen and Llama backbone models. All models are trained with up to five reasoning steps.
@@ 5.2 Effect of Maximum Iteration Steps
In this section, we analyze how the number of iteration steps affects model performance. As shown in Figure 3, R3-RAG demonstrates steady improvements in answer accuracy as the number of steps increases from 1 to 10. Notably, R3-RAG continues to benefit from additional reasoning steps at inference time, even beyond the maximum number used during training (five steps). This suggests that our framework enables models to generalize their reasoning procedures effectively to longer inference trajectories. Furthermore, we observe that R3-RAG consistently outperforms R3-RAG-CS when evaluated with the same maximum number of steps, highlighting the effectiveness of our RL training approach in enabling models to more accurately answer questions while utilizing an equivalent number of reasoning iterations.
@@ 5.3 Impact of Retrieval Top-k Parameter
Figure 4: R3-RAG performance across different retrieval top-k.
We analyze how the number of documents retrieved at each step affects our method’s performance. Figure 4 demonstrates that increasing top-k from 1 to 5 significantly improves performance by approximately 8%, while adding more documents beyond 5 yields minimal gains (less than 1%). Additionally, we observe that our RL approach substantially enhances answer accuracy across all top-k configurations, demonstrating the robustness of our RL method regardless of retrieval volume.
Methods Retriever Hotpot 2Wiki Avg
Llama-3.1-8B
@@
Table 3: Performance comparison of different retrieval methods (BM25, E5, BGE) across models on HotpotQA and 2WikiMultiHopQA datasets. Results show accuracy (ACC) metric only.
5.4 Transferability over Different Retrievers
We analyze the impact of different retrievers on our method compared with several baselines and show the results in Table 3. Despite being trained exclusively with the E5 retriever, R3-RAG maintains the performance and consistently outperforms all baselines when using different retrievers, demonstrating the transferability of R3-RAG’s learned reasoning and retrieval strategies across external retrieval environment. Meanwhile, R3-RAG’s performance variance across three different retrievers is lower than 3%, which further shows R3-RAG’s robustness. In contrast, baseline methods show significant fluctuation. For example, IRCoT’s accuracy on the 2Wiki dataset plummets from 35.8% when using E5 to just 20.1% with BGE, revealing the human-designed workflow’s sensitivity to different external retrieval environment. Additional results are provided in Appendix F.
@@ 5.5 Case Study
Figure 5: Case Study: R3-RAG can reason and retrieve step by step, and adaptively modify the query when it meets retrieval failure.
We conduct a case study on a multi-hop question, as shown in Figure 5, which asks for the film, The Model Couple or Thendral Veesum, whose director was born earlier. RAG + CoT fails to retrieve the comprehensive information. Specifically, it only retrieves who are these movies’ directors and thus fails to generate the answer. In contrast, R3-RAG can decompose the whole question into the following chain of retrieval queries: 1) the director of the film The Model Couple -> William Klein; 2) the director of the film Thendral Veesum -> B. S. Ranga; 3) the birth day of William Klein 4) the birth day of B. S. Ranga. By iterative reasoning and retrieval, R3-RAG can retrieve relevant information progressively and thus get the comprehensive information of the given question. Additionally, R3-RAG can adaptively decompose the previous query when it fails to retrieve relevant information. Specifically, then it tries to retrieve the documents which contain both directors’ birth day and fails, it can further decompose this query into two separate queries and then successfully gets the corresponding information, which shows R3-RAG’s potential when it meets complicated queries.

### v2

@@ 3.3 Reinforcement Learning
Although the cold start can enable the model to learn the manner of iterative reasoning and retrieval, it is hard to make the model fully explore the external retrieval environment, thus leaving it limited by the cold-start data. To address these challenges, we use RL to further harness the model’s own abilities to better reason and retrieve external information. Specifically, we propose two rewards for R3-RAG: 1) answer correctness as an outcome reward, which judges whether the trajectory leads to correct answers by ground truth; and 2) relevance-based document verification as a process reward, encouraging the model to retrieve documents that are highly relevant to users’ questions. In this section, we first introduce the reward design and then the training objective.
3.3.1 Reward Design
@@
The high-quality outcome reward is critical for the model to generate the correct trajectory (DeepSeek-AI et al., 2025). For question-answering, string matching is usually used to judge whether the prediction is correct according to the ground-truth answer. However, since the same entity can have different variants, the match-based judgment can be inaccurate, which introduces extra noise to the training signal. We propose to combine the match-based judgment and the model-based judgment as follows:
@@
where $a$ is the predicted answer, with $\mathrm{Acc}_{\text{match}}$ and $\mathrm{Acc}_{\text{model}}$ denoting the match-based and model-based answer judgments, respectively.
@@
Using only the outcome reward is challenging because it provides a limited fine-grained training signal and results in sparse rewards, which degrades RL effectiveness. We further propose relevance-based document verification to give fine-grained guidance to the process of iterative retrieval. In each retrieval step, we calculate the document relevance as:
$\mathrm{Rel(d_{i}^{j})}=\mathrm{LLM}(I_{relevance},q_{i},d_{i}^{j}),$ (3)
where $I_{relevance}$ is the instruction for the LLM to judge the relevance between the question $q_{i}$ and the document $d_{i}^{j}$, with scores ranging from $0,0.1,0.2,\dots$ to $1$. Since $q_{i}$ and $d_{i}^{j}$ are concatenated and then sent to the LLM, the relevance judgment can take advantage of full token-level interaction and thus guide the model to better retrieve relevant information.
@@
We also implement a format reward $Val(s_{i}^{j})$ to encourage the model to response in the required format as follows:
$\mathrm{Val(s_{i}^{j})}=\begin{cases}1&\text{if }s_{i}^{j}\text{ is valid}\\ 0&otherwise\end{cases}.$ (4)
@@
The overall reward $r(s_{i}^{j})$ is defined as:
$r(s_{i}^{j})=Val(s_{i}^{j})\cdot(\mathrm{Acc}(a_{i})+\mathrm{Rel}(d_{i}^{j}))+Val(s_{i}^{j})-1,$ (5)
where if $s_{i}^{j}$ denotes the final step, $a_{i}$ represents the predicted answer, and $\mathrm{Rel}(d_{i}^{j})$ is set to 0 since no retrieved document $d_{i}^{j}$ is involved in this step. Otherwise, if $s_{i}^{j}$ corresponds to an intermediate step, $d_{i}^{j}$ indicates the retrieved documents at that step, and $\mathrm{Acc}(a_{i})$ is set to 0 as no answer is available.
In this way, we can teach the model to iteratively invoke the retriever and obtain external information to achieve the correct final answer. To enhance the effect of answer-correctness-based reward and prevent the training from being distracted by other rewards, we use the following factor to adjust the process reward throughout the trajectory:
$r_{all}(s_{i}^{j})=r(s_{i}^{j})\cdot\lambda_{T}(T_{i}),$ (6)
where $\lambda_{T}(T_{i})$ is determined based on the trajectory. For trajectories with correct answers, set $\lambda_{T}(T_{i})>1$ to reinforce successful reasoning; for incorrect answers, set $0<\lambda_{T}(T_{i})<1$ to down-weight the process reward; for trajectories with format errors, set $\lambda_{T}(T_{i})=1$ to isolate format penalties to specific steps.
@@
Table 1: Performance comparison of different models and retrieval methods across multi-hop QA datasets. Our R3-RAG method consistently outperforms all baselines across accuracy (ACC), F1, and exact match (EM) metrics, regardless of the base model or retriever used. Bold values indicate the best performance in each column and model group. Missing values (-) indicate results not reported in the original papers.
3.3.2 Training Objective
Guided by the outcome and process reward, we apply the Proximal Policy Optimization (PPO) to train R3-RAG-CS model:
$\mathcal{L}_{RL}=$ $\mathbb{E}_{(x,y)\sim\mathcal{D}_{\mathrm{RL}}^{\mathrm{train}}}\Big[\min\Big(\rho(\theta)A(x,y),$
$\operatorname{clip}(\rho(\theta),1-\epsilon,1+\epsilon)A(x,y)\Big)\Big]$ (7)
$+\beta\mathcal{L}_{KL},$
where $\rho(\theta)=\frac{P(y|x;\theta)}{P(y|x;\theta_{old})}$ represents the probability ratio between the current policy and the old policy, $A(x,y)$ denotes the advantage function estimated using Generalized Advantage Estimation (GAE), $\epsilon$ is the clipping parameter that constrains policy updates, $\mathcal{L}_{KL}$ represents the Kullback-Leibler (KL) divergence between the current policy and the reference model to prevent excessive deviation, and $\beta$ controls the strength of the KL penalty (Schulman et al., 2017). Details are provided in Appendix B.
4 Experiments
@@ 4.1 Datasets and Metrics
We follow Song et al. (2025) to conduct a comparison on three multi-hop QA datasets: HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022), using their full dev sets. In the RL stage, we use only 8,192 examples from HotpotQA for training. We evaluate using three complementary metrics: Exact Match (EM) measures strict string matching against ground-truth answers; F1 score quantifies partial matching via token overlap between predicted and ground-truth answers; and Accuracy (Acc) is based on LLM judgment of semantic correctness. For EM and F1, we first extract the short answer from the LLM’s response and then compare it against the ground-truth answer.
@@ 4.2 Baselines
We compare against several competitive baselines: 1) CoT: LLMs directly answer the question using CoT; 2) RAG with CoT: LLMs answer the question based on documents retrieved using original questions as queries; 3) Iterative retrieval methods: LLMs are prompted to invoke retrieval tools during reasoning, including ReAct Yao et al. (2023), FLARE Jiang et al. (2023), Self-Ask Press et al. (2023), ITER-RETGEN Shao et al. (2023), and IRCoT Trivedi et al. (2023); and 4) Fine-tuned methods, including Auto-RAG Yu et al. (2024), DeepRAG Guan et al. (2025), and ReSearch Chen et al. (2025).
@@ 4.3 Implementation Details
For the cold-start stage, we create 51,254 solution trajectories from the training sets of HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022). For the RL stage, we sample 8,192 instances from the remaining HotpotQA training data. We employ E5-base-v2 (Wang et al., 2022) as our default retriever. During evaluation, the models receive complete retrieved documents, limited to the top 5 per iteration with a maximum of 5 iterations. Unlike prior work (Zheng et al., 2025), which extracts information from retrieved documents to reduce context length, our approach directly incorporates complete retrieved documents. All methods are evaluated using the same metrics and hyperparameters. Further details are provided in Appendix C.
@@ 4.4 Main Results
We show the evaluation results of R3-RAG using Llama-3.1-8B and Qwen2.5-7B in Table 1. We see that R3-RAG significantly outperforms baselines on three datasets across different metrics, which indicates strong overall reasoning and retrieval performance. Specifically, R3-RAG outperforms the strongest iterative RAG workflow method, IRCoT, by around 15 percentage points on average, which shows that R3-RAG can help the model better learn to iteratively reason and retrieve step-by-step and thus get more accurate answers. Additionally, R3-RAG also outperforms ReSearch (Chen et al., 2025), which uses RL with an outcome-based reward to augment the model’s iterative retrieval ability. This shows the advantages of R3-RAG’s reward design, which not only teaches the model to get the final answer but also guides the model to retrieve documents relevant to the question. Meanwhile, R3-RAG achieves significant improvements with both Qwen2.5-7B and Llama-3.1-8B, and this shows the robustness of R3-RAG. Although R3-RAG’s RL is only trained on the HotpotQA dataset, R3-RAG outperforms the cold-start model on three datasets, which shows that the proposed RL algorithm can generalize well across various datasets and further demonstrates R3-RAG’s transferability to different downstream scenarios.
@@ 5.1 Effect of Different Reward
Table 2 presents the comparative performance of different reinforcement learning configurations in R3-RAG. Our full model employs both process and outcome rewards. When we remove the process reward component from our complete R3-RAG framework, we observe a performance decline of 1.4 percentage points on average (from 46.6% to 45.2%), which shows that fine-grained process rewards effectively guide the model to retrieve more relevant documents at each reasoning step, leading to improved overall performance. Further removing the outcome reward component (thus eliminating all reinforcement learning signals) results in an additional performance drop of 3.6 percentage points on average (from 45.2% to 41.6%). These results demonstrate the effectiveness of both reward components. Outcome rewards guide the model to optimize for correct final answers, while process rewards provide granular feedback on the quality of each retrieval operation. By jointly optimizing for both the ultimate goal (correct answers) and the intermediate steps (relevant document retrieval), our approach achieves a more effective reasoning-retrieval strategy.
[…]
@@
Figure 3: Impact of the maximum number of reasoning steps on HotpotQA and 2WikiMultiHopQA. Results are shown for both Qwen and Llama backbone models. All models are trained with up to five reasoning steps.
@@ 5.2 Effect of Maximum Iteration Steps
In this section, we analyze how the number of reasoning steps affects model performance. As shown in Figure 3, R3-RAG demonstrates steady improvements in answer accuracy as the number of steps increases from 1 to 10. Notably, R3-RAG continues to benefit from additional reasoning steps at inference time, even beyond the maximum number used during training (five steps). This suggests that our framework enables models to generalize their reasoning procedures effectively to longer inference trajectories. Furthermore, we observe that R3-RAG consistently outperforms R3-RAG-CS when evaluated with the same maximum number of steps, highlighting the effectiveness of our RL training approach in enabling models to more accurately answer questions while using the same number of reasoning steps.
@@ 5.3 Impact of Retrieval Top-$k$ Parameter
Figure 4: R3-RAG performance across different retrieval top $k$.
We analyze how the number of documents retrieved at each step affects our method’s performance. Figure 4 demonstrates that increasing top $k$ from 1 to 5 significantly improves performance by approximately 8 percentage points, while adding more documents beyond 5 yields minimal gains (less than 1%). Additionally, we observe that our RL approach substantially enhances answer accuracy across all top-$k$ configurations, demonstrating the robustness of our RL method regardless of variation in the number of retrieved documents per step.
5.4 Transferability over Different Retrievers
[…]
@@ 5.4 Transferability over Different Retrievers
Table 3: Performance comparison of different retrieval methods (BM25, E5, BGE) across models on the HotpotQA and 2WikiMultiHopQA datasets. Results show only the accuracy (ACC) metric .
We analyze the impact of different retrievers on our method compared with several baselines and show the results in Table 3. Despite being trained exclusively with the E5 retriever, R3-RAG maintains its performance and consistently outperforms all baselines when using different retrievers, demonstrating the transferability of R3-RAG’s learned reasoning and retrieval strategies across external retrieval environments. Meanwhile, R3-RAG’s performance variation across three different retrievers is less than 3%, which further shows R3-RAG’s robustness. In contrast, baseline methods show significant fluctuation. For example, IRCoT’s accuracy on the 2WikiMultiHopQA dataset plummets from 35.8% when using E5 to just 20.1% with BGE, revealing the human-designed workflow’s sensitivity to different external retrieval environments. Additional results across more datasets and metrics are provided in Appendix D.
5.5 Efficiency Analysis
Method 2Wiki MuSiQue Avg
ReSearch 506.97 541.97 524.46
R3-RAG 382.08 429.94 405.01
Reduction (%) $\downarrow$ 24.6 20.7 22.8
Table 4: Token usage comparison with ReSearch on the Qwen backbone. Values are average LLM tokens per correctly answered question across the multi-step interaction (lower is better).
Longer responses that externalize reasoning often improve accuracy, but they also increase latency and token cost. To quantify this trade-off, we compare R3-RAG with ReSearch (Chen et al., 2025) on the Qwen backbone, reporting the average LLM token usage for correctly answered questions on 2WikiMultiHopQA and MuSiQue. As shown in Table 4, R3-RAG uses 22.8% fewer tokens on average (24.6% on 2WikiMultiHopQA; 20.7% on MuSiQue) while maintaining higher accuracy (Table 1). We attribute this efficiency to our outcome-plus-process rewards: by jointly rewarding the final answer and each intermediate retrieval decision, R3-RAG retrieves more relevant documents and requires fewer, more productive reasoning turns than the outcome-only ReSearch baseline.
@@ 5.6 Case Study
We conduct a case study on a multi-hop question (Figure 5) that asks for the film-The Model Couple or Thendral Veesum-whose director was born earlier. RAG + CoT fails to retrieve comprehensive information for this question. Specifically, it only retrieves the directors of these movies and thus fails to generate the answer. In contrast, R3-RAG can decompose the whole question into the following chain of retrieval queries: 1) the director of the film The Model Couple -> William Klein; 2) the director of the film Thendral Veesum -> B. S. Ranga; 3) the birth date of William Klein; and 4) the birth date of B. S. Ranga. Through iterative reasoning and retrieval, R3-RAG progressively retrieves relevant information and thus obtains comprehensive information for the question. Additionally, R3-RAG can adaptively decompose the previous query when it fails to retrieve relevant information. Specifically, when it tries to retrieve documents that contain both directors’ birth dates and fails, it further decomposes this query into two separate queries and then successfully retrieves the corresponding information, which shows R3-RAG’s potential when it encounters complex queries.
Figure 5: Case Study: R3-RAG can reason and retrieve step-by-step and adaptively modify the query when retrieval fails.